OpenStudy (anonymous):

I have 5 pieces of paper, each with a different math problem. In how many ways can he give these problems to his 10 friends?

OpenStudy (jack1):

hey @JuanSalitras

OpenStudy (jack1):

im not great with permutations or combinations... but I think i can do this one, lemme know if it seems wrong tho and i''ll ask someone else to step in who knows more than i do about them

OpenStudy (jack1):

now, i'm pretty sure this one falls under "permutations": as we can give more than one question to one person, it should be fairly simple: so there are 10 possible choices and 5 questions so 10 x 10 x 10 x 10 x 10 n^r = 10^5 = 100000 ... yeah?

OpenStudy (anonymous):

Thank you so much for your help.

OpenStudy (jack1):

ur right dude, any time

Directrix (directrix):

Why would the answer not be 10*9*8*7*6 ? I am not saying that 10^5 is incorrect but wandering why 10*9*8*7*6 = 30,240 would not be correct. 5 people and 10 problems: first person has 10 problem choices, second has 9, third has 8, fourth has 7, and third has 6.

OpenStudy (jack1):

the OP has edited the question, initially he stated that more than 1 question could go to the same person, therefore the choices are not reduced with each new question ie exponential solve rather than factorial @Directrix

OpenStudy (jack1):

also, u have it backwards, 10 people and 5 problems, not 5 people and 10 problems

Directrix (directrix):

>>the OP has edited the question -- that explains my confusion. Thanks.